# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ 11.5 Learning Rewards from Preferences
The canonical RLHF framework is built on a formal model of human preference. The observed data consists of tuples $(s,y_{w},y_{l})$, where $s$ is a context, and $y_{w}$ and $y_{l}$ are two outputs, with $y_{w}$ being the “winner” preferred by a human. Assuming preferences follow a latent utility model, the Bradley-Terry model (Bradley and Terry, 1952) gives the probability that $y_{w}$ is preferred: $P(y_{w}\succ y_{l}|s)=\sigma(r_{\theta}(s,y_{w})-r_{\theta}(s,y_{l}))$, where $r_{\theta}:\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is a learned reward model parameterized by $\theta$, trained to approximate human preferences (not the ground-truth reward, which is unobserved), and $\sigma(\cdot)$ is the logistic function. This formulation is a binary logit model (Section 2, Equation 1).213213213Iskhakov et al. (2020) discuss the contrasts and synergies between machine learning and structural econometrics, including the shared reliance on logit-based choice models that underlies both RLHF and dynamic discrete choice estimation. The reward model parameters $\theta$ are estimated by minimizing the negative log-likelihood of the observed human choices:
@@
In the LLM setting, the “outputs” $y_{w}$ and $y_{l}$ are token sequences, i.e., trajectories of the autoregressive policy214214214An autoregressive language model generates text token by token: at each step it outputs a distribution over the vocabulary conditioned on all preceding tokens, then samples the next token; the full response is therefore a trajectory in token space and the model acts as a sequential policy over a vocabulary-sized action set., so preferences are over trajectories rather than single actions. The preference loss in Equation (109) is MLE of a choice model where the “alternatives” are trajectory segments and the “choice” is the human-preferred one.
@@ 11.6 The RLHF Pipeline and Direct Optimization
The learned $r_{\theta}$ then serves as a proxy objective for policy optimization. Building on the reward-learning and RL fine-tuning framework of Ziegler et al. (2019) and Stiennon et al. (2020), the canonical three-stage pipeline was formalized by Ouyang et al. (2022). First, a base pretrained model is initialized via supervised fine-tuning (SFT)215215215Pretraining optimizes a language model to predict the next token across a massive text corpus, producing broad linguistic knowledge with no behavioral objective. Supervised fine-tuning (SFT) continues training on a small curated dataset of (prompt, ideal-response) pairs to specialize the model toward the desired task and establish the reference policy $\pi^{SFT}$ from which the KL penalty is measured. on a small dataset of high-quality demonstrations, yielding an initial policy $\pi^{SFT}$. This grounds the model in the desired style and format. The second step is the reward model training as described, using preference data generated from this $\pi^{SFT}$.
In the third step, the SFT policy $\pi_{\phi}$ is fine-tuned via PPO (Section 5.5) to maximize the frozen reward model $r_{\theta}$,216216216After fine-tuning concludes, the resulting LLM is deployed with frozen weights. Each user interaction is a forward pass in the execution phase (Section 2); the model does not update its parameters from conversations. Periodic retraining on new preference data constitutes a separate training phase. with a KL-divergence penalty preventing the policy from drifting into regions where $r_{\theta}$ is unreliable. The objective is
@@ 11.7 Recent Developments
While RLHF and its successors have proven effective at aligning model behavior with human preferences, several open issues remain. The framework does not aim to solve the identification problem in the strict econometric sense; the learned reward model $r_{\theta}$ (whether explicit or implicit in DPO) is a proxy for preference, not necessarily the uniquely identified “true” utility function required for welfare analysis.218218218The identification problem in preference learning mirrors that in discrete choice. The reward function is identified only up to an additive constant and requires a location normalization. Furthermore, the notion of “human feedback” is a significant simplification, as the preferences being optimized are those of a small, non-representative group of paid labelers, raising important questions about whose values are being embedded in these systems. Finally, the fine-tuning process can lead to degraded performance on standard academic benchmarks, a phenomenon called the alignment tax (Ouyang et al., 2022).219219219The alignment tax, as described by Ouyang et al. (2022), refers to performance regressions on public NLP benchmarks (SQuAD, DROP, HellaSwag) that result from RLHF fine-tuning. Mitigating these challenges while scaling alignment beyond the limits of human data collection remains a key frontier for the field.
@@ 11.8 Simulation Study: Preference Learning in Job Search
Six methods are compared across $K\in\{25,50,100,200,500,1{,}000,2{,}000,5{,}000\}$, averaged over 30 seeds. The neural network RLHF reward model is a two-layer MLP trained by Bradley-Terry MLE on the $K$ segment pairs; per-transition rewards are discount-weighted and summed to obtain a segment score, and the resulting 112-state reward table is solved by value iteration.220220220The neural network has 4 inputs (normalized log-wage, normalized amenity, employment indicator, action), 32 hidden units per layer, and $\sim$1,200 parameters. The logistic loss from Equation (109) is applied to discount-weighted segment reward sums. The correctly specified structural model parameterizes utility as $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})z$, estimating the single parameter $\hat{\alpha}$ via Bradley-Terry MLE, then solves the MDP. The misspecified model uses $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})\bar{z}$, where $\bar{z}=3$ is the mean amenity level; this model ignores amenity variation across jobs, treating all amenities as identical. DPO trains a tabular softmax policy directly from trajectory comparisons, bypassing reward modeling entirely.221221221DPO uses 112 logit parameters $\phi_{s}$, one per state, trained via the DPO loss (Equation 113) with Adam optimization over a sweep of $\lambda_{KL}\in\{0.01,0.05,0.1,0.5,1.0,5.0\}$, selecting the $\lambda_{KL}$ that minimizes training loss. The reference policy is uniform: $\pi^{SFT}(a|s)=0.5$. To match the LLM setup where both completions condition on the same prompt, DPO comparison pairs start from the same initial state. Tabular Q-learning (10,00010{,}000 episodes, $\varepsilon=0.15$, learning rate $0.1$) and exact DP provide scalar-reward baselines. All four preference methods receive identical comparison data per seed; DPO receives a same-state variant generated from the same seed.
@@
Figure 24 reports the main results. Three findings stand out. First, the correctly specified structural model dominates: even at $K=25$ it achieves 99.9% of DP-optimal, illustrating the power of correct specification when the model has a single free parameter. Second, the neural network converges more slowly but reaches 99.9% by K=5, [^0] K=5{,}000, reflecting the higher sample complexity of a flexible model relative to a one-parameter structural specification. Third, DPO plateaus at approximately 95% by $K=500$ and does not improve further.222222222DPO learns only from $(s,a)$ pairs visited in training trajectories generated by the random behavioral policy. It cannot propagate value to undervisited states the way value iteration does after learning a reward model, so states poorly covered by the behavioral policy remain suboptimal regardless of $K$. This plateau is qualitatively different from the gridworld failure ($118\%$): DPO recovers a reasonable policy, but the gap does not close with additional data.223223223DPO fails catastrophically in gridworld because transitions are stochastic (10% slip probability) and rewards are transition-dependent. The same $(s,a)$ pair yields different rewards depending on whether the agent slipped, so the DPO loss conflates policy quality with transition luck. In the job search model, accept/reject deterministically changes employment status, and only the 5% layoff probability introduces stochastic transitions. The misspecified constant-amenity model plateaus at 91% regardless of $K$, because additional preference data cannot correct the omitted variable.
@@
Table 21 provides state-level diagnostics at K=5, [^0] K=5{,}000. The structural model and neural network achieve near-perfect policy agreement with $\pi^{*}$ (100% and 96% respectively). DPO agrees on only 57% of states with value-function correlation 0.78, systematically underselecting on both wage and amenity dimensions.224224224DPO’s mean accepted amenity of 3.4 parallels the misspecified structural model’s 3.0, though the mechanisms differ: the misspecified model ignores amenity variation by construction, while DPO underweights amenities because the random behavioral policy underrepresents high-amenity employed states in the training data. The misspecified model agrees on 50%, with disagreements concentrated where amenity variation matters.225225225An online versus offline ablation for the neural network at K=1, [^0] K=1{,}000 (20 seeds) shows comparable performance: online $73.92\pm 0.05$, offline $73.99\pm 0.03$ ($p=0.09$). The random behavioral policy already provides diverse career trajectories covering the full wage-amenity space.
@@
Two-stage RLHF separates preference estimation (a static econometric problem) from dynamic programming (which exploits the known transition model); DPO conflates the two and forfeits the transition structure that structural modelers typically have access to.226226226This advantage is specific to settings where the transition model is known or estimable; in domains without a tractable transition model, DPO’s single-stage approach avoids compounding errors from reward model estimation.
@@ 12 Reinforcement Learning and Causal Inference
Reinforcement learning algorithms solve Markov decision processes by estimating value functions or optimizing policies from sampled transitions. Two assumptions underlie the standard formulation. First, the agent’s action at time $t$ is determined solely by the observed state $s_{t}$ and the agent’s policy $\pi(a\mid s_{t})$; there is no unobserved variable (confounder) simultaneously influencing both the action and the reward or state transition.227227227See Section 2 for the terminological mapping between “outcome” in causal inference and the corresponding RL quantities. Throughout this chapter, I reserve “outcome” for its causal inference meaning and use the specific RL quantity (reward, state, return, value) elsewhere. Second, the observed state $s_{t}$ is sufficient for prediction, so that conditioning on $s_{t}$ renders future states independent of past history. Both assumptions are the Markov property restated in causal language. Both fail routinely in economic applications.
@@ 12.2 The Confounded MDP
When unobserved confounders influence both the behavioral policy and the transitions or rewards, the MDP is confounded. This formalization, developed by Zhang and Bareinboim (2019), Zhang and Bareinboim (2020), and Kallus and Zhou (2020), provides the foundation for causal reasoning in sequential decision problems. The key tool is Pearl’s do-operator. The interventional distribution $P(Y\mid\operatorname{do}(X=x))$ is the distribution that arises when $X$ is set externally rather than observed passively, severing all incoming causal influences on $X$ while leaving the remaining data-generating process intact.228228228The do-operator is formalized within the structural causal model (SCM) framework of Pearl (2009). An SCM specifies endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and a distribution $P(\mathbf{U})$. The intervention $\operatorname{do}(X=x)$ replaces the structural equation for $X$ with a constant, producing the interventional distribution. See Pearl (2009) for the complete framework, including the causal hierarchy (association, intervention, counterfactual) and general identification theory.
@@
The importance-sampling estimator $\hat{V}^{\pi}_{\text{IS}}=\frac{1}{N}\sum_{i=1}^{N}\prod_{t=0}^{T}\frac{\pi(a_{t}^{(i)}\mid s_{t}^{(i)})}{\mu(a_{t}^{(i)}\mid s_{t}^{(i)})}G^{(i)}$, where $G^{(i)}=\sum_{t=0}^{T}\gamma^{t}R_{t}^{(i)}$ is the discounted return of trajectory $i$ and $T$ is the trajectory length, is also biased because the propensity $\mu(a\mid s)$ is not the true behavioral propensity $\mu(a\mid s,u)$.229229229This is the sequential analogue of the omitted variable bias in linear regression. In the static case, regressing $Y$ on $X$ without controlling for a confounder $U$ yields a biased coefficient. In the sequential case, the bias propagates through the Bellman recursion and can amplify over the horizon.
@@
Suppose a set of observed variables $\mathbf{Z}_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$ in the causal graph of the confounded MDP, meaning that $\mathbf{Z}_{t}$ blocks all backdoor paths from $A_{t}$ to $S_{t+1}$ and no element of $\mathbf{Z}_{t}$ is a descendant of $A_{t}$.230230230A backdoor path from $A_{t}$ to $S_{t+1}$ is any path in the causal graph that begins with an arrow into $A_{t}$ (i.e., a non-causal path). In the confounded MDP, $A_{t}\leftarrow U_{t}\rightarrow S_{t+1}$ is a backdoor path: $U_{t}$ causes both $A_{t}$ and $S_{t+1}$, creating a spurious association. Blocking all such paths by conditioning on appropriate variables eliminates the confounding bias. See Pearl (2009), Chapter 3. Then the interventional transition probability is identified:
@@ 12.3 Backdoor-Adjusted Off-Policy Evaluation
The doubly robust variant combines the fitted action-value function $\hat{Q}(s,a)$ (the RL analogue of the econometric outcome model) with backdoor-adjusted propensities, achieving consistency if either $\hat{Q}$ or the propensity model is correctly specified.

### 24 Mar 2026 v5

@@ 11.5 Learning Rewards from Preferences
The canonical RLHF framework is built on a formal model of human preference. The observed data consists of tuples $(s,y_{w},y_{l})$, where $s$ is a context, and $y_{w}$ and $y_{l}$ are two outputs, with $y_{w}$ being the “winner” preferred by a human. Assuming preferences follow a latent utility model, the Bradley-Terry model (Bradley and Terry, 1952) gives the probability that $y_{w}$ is preferred: $P(y_{w}\succ y_{l}|s)=\sigma(r_{\theta}(s,y_{w})-r_{\theta}(s,y_{l}))$, where $r_{\theta}:\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is a learned reward model parameterized by $\theta$, trained to approximate human preferences (not the ground-truth reward, which is unobserved), and $\sigma(\cdot)$ is the logistic function. This formulation is a binary logit model (Section 2, Equation 1).215215215Iskhakov et al. (2020) discuss the contrasts and synergies between machine learning and structural econometrics, including the shared reliance on logit-based choice models that underlies both RLHF and dynamic discrete choice estimation. The reward model parameters $\theta$ are estimated by minimizing the negative log-likelihood of the observed human choices:
@@
In the LLM setting, the “outputs” $y_{w}$ and $y_{l}$ are token sequences, i.e., trajectories of the autoregressive policy216216216An autoregressive language model generates text token by token: at each step it outputs a distribution over the vocabulary conditioned on all preceding tokens, then samples the next token; the full response is therefore a trajectory in token space and the model acts as a sequential policy over a vocabulary-sized action set., so preferences are over trajectories rather than single actions. The preference loss in Equation (109) is MLE of a choice model where the “alternatives” are trajectory segments and the “choice” is the human-preferred one.
@@ 11.6 The RLHF Pipeline and Direct Optimization
The learned $r_{\theta}$ then serves as a proxy objective for policy optimization. Building on the reward-learning and RL fine-tuning framework of Ziegler et al. (2019) and Stiennon et al. (2020), the canonical three-stage pipeline was formalized by Ouyang et al. (2022). First, a base pretrained model is initialized via supervised fine-tuning (SFT)217217217Pretraining optimizes a language model to predict the next token across a massive text corpus, producing broad linguistic knowledge with no behavioral objective. Supervised fine-tuning (SFT) continues training on a small curated dataset of (prompt, ideal-response) pairs to specialize the model toward the desired task and establish the reference policy $\pi^{SFT}$ from which the KL penalty is measured. on a small dataset of high-quality demonstrations, yielding an initial policy $\pi^{SFT}$. This grounds the model in the desired style and format. The second step is the reward model training as described, using preference data generated from this $\pi^{SFT}$.
In the third step, the SFT policy $\pi_{\phi}$ is fine-tuned via PPO (Section 5.5) to maximize the frozen reward model $r_{\theta}$,218218218After fine-tuning concludes, the resulting LLM is deployed with frozen weights. Each user interaction is a forward pass in the execution phase (Section 2); the model does not update its parameters from conversations. Periodic retraining on new preference data constitutes a separate training phase. with a KL-divergence penalty preventing the policy from drifting into regions where $r_{\theta}$ is unreliable. The objective is
@@ 11.7 Recent Developments
While RLHF and its successors have proven effective at aligning model behavior with human preferences, several open issues remain. The framework does not aim to solve the identification problem in the strict econometric sense; the learned reward model $r_{\theta}$ (whether explicit or implicit in DPO) is a proxy for preference, not necessarily the uniquely identified “true” utility function required for welfare analysis.220220220The identification problem in preference learning mirrors that in discrete choice. The reward function is identified only up to an additive constant and requires a location normalization. Furthermore, the notion of “human feedback” is a significant simplification, as the preferences being optimized are those of a small, non-representative group of paid labelers, raising important questions about whose values are being embedded in these systems. Finally, the fine-tuning process can lead to degraded performance on standard academic benchmarks, a phenomenon called the alignment tax (Ouyang et al., 2022).221221221The alignment tax, as described by Ouyang et al. (2022), refers to performance regressions on public NLP benchmarks (SQuAD, DROP, HellaSwag) that result from RLHF fine-tuning. Mitigating these challenges while scaling alignment beyond the limits of human data collection remains a key frontier for the field.
@@ 11.8 Simulation Study: Preference Learning in Job Search
Six methods are compared across $K\in\{25,50,100,200,500,1{,}000,2{,}000,5{,}000\}$, averaged over 30 seeds. The neural network RLHF reward model is a two-layer MLP trained by Bradley-Terry MLE on the $K$ segment pairs; per-transition rewards are discount-weighted and summed to obtain a segment score, and the resulting 112-state reward table is solved by value iteration.222222222The neural network has 4 inputs (normalized log-wage, normalized amenity, employment indicator, action), 32 hidden units per layer, and $\sim$1,200 parameters. The logistic loss from Equation (109) is applied to discount-weighted segment reward sums. The correctly specified structural model parameterizes utility as $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})z$, estimating the single parameter $\hat{\alpha}$ via Bradley-Terry MLE, then solves the MDP. The misspecified model uses $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})\bar{z}$, where $\bar{z}=3$ is the mean amenity level; this model ignores amenity variation across jobs, treating all amenities as identical. DPO trains a tabular softmax policy directly from trajectory comparisons, bypassing reward modeling entirely.223223223DPO uses 112 logit parameters $\phi_{s}$, one per state, trained via the DPO loss (Equation 113) with Adam optimization over a sweep of $\lambda_{KL}\in\{0.01,0.05,0.1,0.5,1.0,5.0\}$, selecting the $\lambda_{KL}$ that minimizes training loss. The reference policy is uniform: $\pi^{SFT}(a|s)=0.5$. To match the LLM setup where both completions condition on the same prompt, DPO comparison pairs start from the same initial state. Tabular Q-learning (10,00010{,}000 episodes, $\varepsilon=0.15$, learning rate $0.1$) and exact DP provide scalar-reward baselines. All four preference methods receive identical comparison data per seed; DPO receives a same-state variant generated from the same seed.
@@
Figure 24 reports the main results. Three findings stand out. First, the correctly specified structural model dominates: even at $K=25$ it achieves 99.9% of DP-optimal, illustrating the power of correct specification when the model has a single free parameter. Second, the neural network converges more slowly but reaches 99.9% by K=5, [^0] K=5{,}000, reflecting the higher sample complexity of a flexible model relative to a one-parameter structural specification. Third, DPO plateaus at approximately 95% by $K=500$ and does not improve further.224224224DPO learns only from $(s,a)$ pairs visited in training trajectories generated by the random behavioral policy. It cannot propagate value to undervisited states the way value iteration does after learning a reward model, so states poorly covered by the behavioral policy remain suboptimal regardless of $K$. This plateau is qualitatively different from the gridworld failure ($118\%$): DPO recovers a reasonable policy, but the gap does not close with additional data.225225225DPO fails catastrophically in gridworld because transitions are stochastic (10% slip probability) and rewards are transition-dependent. The same $(s,a)$ pair yields different rewards depending on whether the agent slipped, so the DPO loss conflates policy quality with transition luck. In the job search model, accept/reject deterministically changes employment status, and only the 5% layoff probability introduces stochastic transitions. The misspecified constant-amenity model plateaus at 91% regardless of $K$, because additional preference data cannot correct the omitted variable.
@@
Table 21 provides state-level diagnostics at K=5, [^0] K=5{,}000. The structural model and neural network achieve near-perfect policy agreement with $\pi^{*}$ (100% and 96% respectively). DPO agrees on only 57% of states with value-function correlation 0.78, systematically underselecting on both wage and amenity dimensions.226226226DPO’s mean accepted amenity of 3.4 parallels the misspecified structural model’s 3.0, though the mechanisms differ: the misspecified model ignores amenity variation by construction, while DPO underweights amenities because the random behavioral policy underrepresents high-amenity employed states in the training data. The misspecified model agrees on 50%, with disagreements concentrated where amenity variation matters.227227227An online versus offline ablation for the neural network at K=1, [^0] K=1{,}000 (20 seeds) shows comparable performance: online $73.92\pm 0.05$, offline $73.99\pm 0.03$ ($p=0.09$). The random behavioral policy already provides diverse career trajectories covering the full wage-amenity space.
@@
Two-stage RLHF separates preference estimation (a static econometric problem) from dynamic programming (which exploits the known transition model); DPO conflates the two and forfeits the transition structure that structural modelers typically have access to.228228228This advantage is specific to settings where the transition model is known or estimable; in domains without a tractable transition model, DPO’s single-stage approach avoids compounding errors from reward model estimation.
@@ 12 Reinforcement Learning and Causal Inference
Reinforcement learning algorithms solve Markov decision processes by estimating value functions or optimizing policies from sampled transitions. Two assumptions underlie the standard formulation. First, the agent’s action at time $t$ is determined solely by the observed state $s_{t}$ and the agent’s policy $\pi(a\mid s_{t})$; there is no unobserved variable (confounder) simultaneously influencing both the action and the reward or state transition.229229229See Section 2 for the terminological mapping between “outcome” in causal inference and the corresponding RL quantities. Throughout this chapter, I reserve “outcome” for its causal inference meaning and use the specific RL quantity (reward, state, return, value) elsewhere. Second, the observed state $s_{t}$ is sufficient for prediction, so that conditioning on $s_{t}$ renders future states independent of past history. Both assumptions are the Markov property restated in causal language. Both fail routinely in applied settings.
@@ 12.2 The Confounded MDP
When unobserved confounders influence both the behavioral policy and the transitions or rewards, the MDP is confounded. This formalization, developed by Zhang and Bareinboim (2019), Zhang and Bareinboim (2020), and Kallus and Zhou (2020), provides the foundation for causal reasoning in sequential decision problems. The key tool is Pearl’s do-operator. The interventional distribution $P(Y\mid\operatorname{do}(X=x))$ is the distribution that arises when $X$ is set externally rather than observed passively, severing all incoming causal influences on $X$ while leaving the remaining data-generating process intact.230230230The do-operator is formalized within the structural causal model (SCM) framework of Pearl (2009). An SCM specifies endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and a distribution $P(\mathbf{U})$. The intervention $\operatorname{do}(X=x)$ replaces the structural equation for $X$ with a constant, producing the interventional distribution. See Pearl (2009) for the complete framework, including the causal hierarchy (association, intervention, counterfactual) and general identification theory.
@@
The importance-sampling estimator $\hat{V}^{\pi}_{\text{IS}}=\frac{1}{N}\sum_{i=1}^{N}\prod_{t=0}^{T}\frac{\pi(a_{t}^{(i)}\mid s_{t}^{(i)})}{\mu(a_{t}^{(i)}\mid s_{t}^{(i)})}G^{(i)}$, where $G^{(i)}=\sum_{t=0}^{T}\gamma^{t}R_{t}^{(i)}$ is the discounted return of trajectory $i$ and $T$ is the trajectory length, is also biased because the propensity $\mu(a\mid s)$ is not the true behavioral propensity $\mu(a\mid s,u)$.231231231This is the sequential analogue of the omitted variable bias in linear regression. In the static case, regressing $Y$ on $X$ without controlling for a confounder $U$ yields a biased coefficient. In the sequential case, the bias propagates through the Bellman recursion and can amplify over the horizon.
@@
Suppose a set of observed variables $\mathbf{Z}_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$ in the causal graph of the confounded MDP, meaning that $\mathbf{Z}_{t}$ blocks all backdoor paths from $A_{t}$ to $S_{t+1}$ and no element of $\mathbf{Z}_{t}$ is a descendant of $A_{t}$.232232232A backdoor path from $A_{t}$ to $S_{t+1}$ is any path in the causal graph that begins with an arrow into $A_{t}$ (i.e., a non-causal path). In the confounded MDP, $A_{t}\leftarrow U_{t}\rightarrow S_{t+1}$ is a backdoor path: $U_{t}$ causes both $A_{t}$ and $S_{t+1}$, creating a spurious association. Blocking all such paths by conditioning on appropriate variables eliminates the confounding bias. See Pearl (2009), Chapter 3. Then the interventional transition probability is identified:
@@ 12.3 Backdoor-Adjusted Off-Policy Evaluation
The doubly robust variant combines the fitted action-value function $\hat{Q}(s,a)$ with backdoor-adjusted propensities, achieving consistency if either $\hat{Q}$ or the propensity model is correctly specified.